1.    y''' - 9y''+ 27y'-27y=0        ans, y=c1e3x+c2xe3x+c3x2e3x

2.    y''' + 2y''- 13y' + 10y =0    ans, y= c1e-5x+c2e2x+c3ex

1
Expert's answer
2021-12-08T10:19:25-0500

 Q1)

y''' - 9y'' + 27y' - 27y = 0

Let us substitute y = emx .

y' = memx , y'' = m²emx and y''' = m³emx

So the auxiliary equation is

m³ - 9m² + 27m - 27 = 0

=> m³ - 3m².3 + 3m.3² - 3³=0

=> (m-3)³=0

=> m = 3,3,3

So the general solution of the differential equation is y = c1e3x + c2xe3x + c3x2e3x

Q2)

y''' + 2y''-13y'+10y = 0

Let us substitute y = emx .

y' = memx , y'' = m²emx and y''' = m³emx

So the auxiliary equation is

So m³ + 2m² - 13m + 10 = 0

=> m²(m-1)+3m(m-1)-10(m-1)=0

=> (m-1)(m²+3m-10)=0

=> (m-1)(m²+5m-2m-10)=0

=> (m-1){m(m+5)-2(m+5)}=0

=> (m-1)(m+5)(m-2)=0

=> m = -5, 2, 1

So the general solution is

y = c1e-5x + c2e2x + c3ex